Can you give me the equation to this word problem. A discount store sold plastic cups for $3.50 each and ceramic cups for $4 each. If 400 cups were sold for a total of $1,458, how many cups of each type were sold? What was the dollar value of each type of cup sold?
Answer by Alan3354(31518)

Can you give me the equation to this word problem. A discount store sold plastic cups for $3.50 each and ceramic cups for $4 each. If 400 cups were sold for a total of $1,458, how many cups of each type were sold? What was the dollar value of each type of cup sold?
----------------
Yes, I can.
P + C = 400 (400 total)
3.5P + 4C = 1458 (total cost)
How's that?